OpenStudy (anonymous):

Can someone show me to get the solution to y''+y'=e^(-x) ?? I have that y_c is r=0 and r=-1 but I can't figure out y_p. Any help would be greatly appreciated.

zepdrix (zepdrix):

\[\Large\rm y''+y'=e^{-x}\]So your homogeneous part of your solution is given by the characteristic equation:\[\Large\rm r^2+r=0\]Ok your roots look good. \[\Large\rm y_c=e^{0x}+e^{-x}\]

zepdrix (zepdrix):

Our particular solution will match the form of the right side,\[\Large\rm y_p=Ae^{-x}\]We just need to know what this A value is.

OpenStudy (anonymous):

I got that part as well. I am using the form of Ae^(-x) for the y_p and have made it down to Ae^(-x) - Ae^(-x) = e^(-x), but I'm lost after this.

OpenStudy (anonymous):

I can see by observation that A = 1, but I'm just not quite sure how to get the finishing touches on the answer that Wolfram Alpha has.

zepdrix (zepdrix):

Hmm yah I was looking at their solution also :o hmmm

zepdrix (zepdrix):

Well we're getting A-A=1, yes? I'm not sure how we get A=1 from that D: Hmm

OpenStudy (anonymous):

oh wait..you are right! yes A is not 1.

zepdrix (zepdrix):

Mmmm, so remember that when you have repeated roots, you multiply by an x each time yah? I guess that's what happening here? I didn't know that could happen with your particular solution :U mmm im rusty...

OpenStudy (anonymous):

yes yes you are completely right. Just as you are saying that I am digging back in my notes and see that! Thank you very much for your help.